OpenStudy (anonymous):

How to rewrite the this as f(x)=mx+b 28x=7y-14

OpenStudy (zzr0ck3r):

add \(14\) to both sides, then divide by \(7\) and tell me what you get

OpenStudy (zzr0ck3r):

or wait for the answer....?

OpenStudy (anonymous):

Well this is the form called y-intercept form, y=mx+b, where ity is being denoted as a function f(x)=mx+b Get y by itself. Add 14 to both sides to get 28x+14=7y... reverse this around to look better 7y = 28x + 14 Divide all terms by 7 to get y = 28/7 + 2

OpenStudy (anonymous):

Replace y with f(x) f(x) = (28/7) + 2

OpenStudy (zzr0ck3r):

I think @Cosmichaotic means \(f(x) = \frac{28}{7}x+2=4x+2\) so \(f(x) = 4x+2\)

OpenStudy (anonymous):

@zzr0ck3r 's suggestion/correction, f(x)=4x+2 is the final form, correct.

OpenStudy (anonymous):

can you give me proper procedure how to put that 14 >.<

OpenStudy (anonymous):

\[28x=7y-14\]\[28x+(-28x)+7y=-28x-14\]\[\frac{ 7y }{ 7}=\frac{ -28-14 }{ 7}\] y=-4x-2